# Rethinking Text-based Protein Understanding: Retrieval or LLM?

### v1

@@ 1 Introduction
  • We perform detailed evaluations on existing protein-text benchmarks, exposing issues related to data leakage and metric deficiencies.
@@ 2.3 Protein Related Retrieval-Based Methods
Retrieval-based approaches are fundamental to protein science, grounded in the well-established biological principle that sequence homology implies evolutionary conservation and functional similarity Pearson (2013). Single-sequence alignment approaches Altschul et al. (1990); Buchfink et al. (2015); Steinegger and Söding (2017); van Kempen et al. (2022) and multiple sequence alignment tools Remmert et al. (2012); Johnson et al. (2010) are extensively used in bioinformatics for identifying highly homologous sequences. Many protein models utilize retrieval methods to assist downstream tasks, where AlphaFold2 Jumper et al. (2021), MSA-Transformer Rao et al. (2021), and RosettaFold Baek et al. (2021) employ multiple sequence alignment results to aid property prediction or structure folding. Furthermore, retrieval-based approaches Tan et al. (2024); Shaw et al. (2024); Sgarbossa and Bitbol (2025); Jin et al. (2024); Li et al. (2024) have demonstrated the feasibility of using retrieval tools to enhance LLM-based predictions in protein research.
@@ 4.3 Novel Benchmark Proposal
To mitigate data leakage, we restructure the protein-text data from Mol-Instruction through a two-step sequence-based partitioning strategy. (1) Low similarity (Low-Sim) split: Based on MMSeqs2 clustering with an 8:2 class split, mitigating general leakage. (2) Out-of-Distribution (OOD): Filters Low-Sim by removing samples with answers present in the training set, preventing reliance on retrieval. We refer to this new benchmark dataset as the Mol-Instructions-OOD dataset. Detailed OOD dataset construction process can be seen in Appendix A.
@@
Table 3: Performance across 4 protein understanding tasks, each evaluated with E-BLEU(Entity-BLEU) and RG-L(ROUGE-L). "*" means using LoRA Hu et al. (2022) fine-tuning. Bold for best, underline for second best.
@@ 5 Experiments
This section comprehensively evaluates the proposed RAPM with the benchmark and metric proposed in Sec. 4.3. Experimental results demonstrate that retrieval-based methods show significant performance degradation in OOD-benchmarks. In ablation studies, we assess the impact of the number of items retrieved, database indexing methods, and prompt construction approaches for retrieval.
@@ 5.1 Experimental Setup
We test four subtasks in Mol-Instructions-OOD, including “Protein Function”, “Functional Description”, “Domain/Motif” and “Catalytic Activity”, using the standard NLP metrics (ROUGE-L) and our proposed metric, Entity-BLEU ($N=2$). Detailed fine-tuning hyperparameters, retrieval settings, and RAPM prompt can be seen in Appendix C.2 and 7. Note that for a fair comparison between RAPM methods and fine-tuned LLMs, we exclude all extra-training-set data during retrieval to prevent potential data leakage.
5.2 OOD Benchmark Performance
Table 3 summarizes the main results on existing benchmarks for four representative methodological approaches, and we observe the following key results: (1) When evaluated in OOD settings, the RAPM method achieves the highest Entity-BLEU scores, outperforming retrieval-based methods and demonstrating substantial improvements over fine-tuned and task-prompted LLMs. Beyond the superior performance, RAPM requires substantially fewer computational resources than fine-tuned LLMs and demonstrates a stronger capability to handle diverse tasks than retrieval-based methods. (2) When comparing ROUGE-L and Entity-BLEU scores of different methods, we observe a poor correlation between them, particularly for fine-tuned LLMs which have high ROUGE-L scores and low Entity-BLEU scores. As discussed in Sec 3.2, we owe this to the fact that fine-tuned LLMs primarily focus on learning irrelevant response patterns rather than understanding protein sequences.
@@ B.2 Details of Protein Knowledge Database Construction
We integrated annotations from three prominent sources: InterPro, EC-GO, and Mol-Instructions. For InterPro, we selectively used only sequences annotated via Swiss-Prot curation, resulting in a high-quality subset with 573,230 sequences. In the EC-GO database, labels corresponding to the enzyme classification (EC) and gene ontology (GO) were merged into a unified text-based annotation to capture functional and process aspects simultaneously. For Mol-Instructions, only the “meta-data” field is retained as the annotation, disregarding the original class labels, to emphasize naturalistic, descriptive phrasing of protein functions.

### v2

@@ 1 Introduction
  • We evaluate existing protein-text benchmarks, revealing data leakage and metric limitations, and propose the new Prot-Inst-OOD dataset and Bio-Entity BLEU metric.
@@ 2.3 Protein Related Retrieval-Based Methods
Retrieval-based approaches are fundamental to protein science, grounded in the well-established biological principle that sequence homology implies evolutionary conservation and functional similarity Pearson (2013). Single-sequence alignment approaches Altschul et al. (1990); Buchfink et al. (2015); Steinegger and Söding (2017); van Kempen et al. (2022) and multiple sequence alignment tools Remmert et al. (2012); Johnson et al. (2010) are extensively used in bioinformatics for identifying highly homologous sequences. Many protein models utilize retrieval methods to assist downstream tasks, where AlphaFold2 Jumper et al. (2021), MSA-Transformer Rao et al. (2021), and RosettaFold Baek et al. (2021) employ multiple sequence alignment results to aid property prediction or structure folding. Furthermore, retrieval-based approaches Tan et al. (2024); Shaw et al. (2024); Sgarbossa and Bitbol (2025); Jin et al. (2024); Li et al. (2024); Ma et al. (2024) have demonstrated the feasibility of using retrieval tools to enhance LLM-based predictions in protein research.
@@ 4.3 Novel Benchmark Proposal
To address data leakage, we reconstruct protein-text datasets by integrating four protein understanding tasks in Mol-Instructions Fang et al. (2024) and captioning tasks in Swiss-Prot Bairoch and Apweiler (2000). For captioning tasks, we generate instructions from original annotations, forming a unified dataset (Pro-Inst-OOD). The OOD construction involves two steps: (1) Low similarity (Low-Sim) split: Based on MMSeqs2 clustering with an 8:2 class split, mitigating general leakage. (2) Out-of-Distribution (OOD): Filters Low-Sim split by removing test samples with answers can be retrieved from the training set, preventing reliance on retrieval. This creates the Pro-Inst-OOD benchmark (construction details in Appendix A).
@@
Table 3: Performance of different approaches in Prot-Inst-OOD, each evaluated with E-BLEU(Entity-BLEU) and RG-L(ROUGE-L). "*" means using LoRA Hu et al. (2022) fine-tuning. Bold for best, underline for second best.
@@ 5 Experiments
This section comprehensively evaluates the proposed RAPM with the benchmark and metric proposed in Sec. 4.3. Experimental results demonstrate that retrieval-based methods show significant performance degradation in the Pro-Inst-OOD benchmark. In ablation studies, we assess the impact of the number of items retrieved, database indexing methods, and prompt construction approaches for retrieval.
@@ 5.1 Experimental Setup
We test all subtasks in Pro-Inst-OOD, including “Protein Function”, “Functional Description”, “Domain/Motif”, “Catalytic Activity”, and “Protein Caption”, using the standard NLP metrics (ROUGE-L) and our proposed metric, Entity-BLEU ($N=2$). Detailed fine-tuning hyperparameters, retrieval settings, and RAPM prompt can be seen in Appendix C.2 and 7. Note that for a fair comparison between RAPM methods and fine-tuned LLMs, we exclude all extra-training-set data during retrieval to prevent potential data leakage. In addition, all subtasks are trained and tested together, but the results are reported separately.
5.2 Pro-Inst-OOD Performance
Table 3 summarizes the main results on existing benchmarks for four representative methodological approaches, and we observe the following key results: (1) When evaluated in OOD settings, the RAPM method achieves the highest Entity-BLEU scores, outperforming retrieval-based methods and demonstrating substantial improvements over fine-tuned and task-prompted LLMs. Beyond the superior performance, RAPM requires substantially fewer computational resources than fine-tuned LLMs and demonstrates a stronger capability to handle diverse tasks than retrieval-based methods. (2) When comparing ROUGE-L and Entity-BLEU scores of different methods, we observe a poor correlation between them, particularly for fine-tuned LLMs, which have high ROUGE-L scores and low Entity-BLEU scores. As discussed in Sec 3.2, we owe this to the fact that fine-tuned LLMs primarily focus on learning irrelevant response patterns rather than understanding protein sequences.
@@ B.2 Details of Protein Knowledge Database Construction
We integrated annotations from three prominent sources: PEERXu et al. (2022), InterProBlum et al. (2025), EC-GOGligorijević et al. (2021), and Mol-InstructionsFang et al. (2024). For InterPro, we selectively used only sequences annotated via Swiss-Prot curation, resulting in a high-quality subset with 573,230 sequences. In the EC-GO database, labels corresponding to the enzyme classification (EC) and gene ontology (GO) were merged into a unified text-based annotation to capture functional and process aspects simultaneously. For Mol-Instructions, only the “meta-data” field is retained as the annotation, disregarding the original class labels, to emphasize naturalistic, descriptive phrasing of protein functions.